f(x)= 2x^2-x+4 (f(x+h)-f(x))/h

Answer by stanbon(75887)

f(x)= 2x^2-x+4 (f(x+h)-f(x))/h
------------------------
f(x+h) = 2(x+h)^2-(x+h)+4
f(x+h) = 2x^2+4hx+2h^2-x-h+4
f(x+h) = 2x^2+4hx-x+2h^2-h+4
-----------
f(x)= 2x^2-x+4
---------
Subtract f(x) from f(x+h) to get:
= 4hx+2h^2-h
--------
Divide by h to get:
[f(x+h)-f(x)]/h = 4x+2h-1
==============
Cheers,
Stan H.